# Additive Complementary Pairs of Codes

### v2

@@ 3 Characterization of ACP of codes
For two ${\rm I\!F}_{q}$-linear additive codes over ${\rm I\!F}_{q^{m}}$, $C$ and $D$, the pair $(C,D)$ is called an additive complementary pair (ACP) of codes if $C\oplus_{{\rm I\!F}_{q}}D={\rm I\!F}_{q^{m}}^{n}$. Equivalently, a pair $(C,D)$ is an ACP of codes if and only if $C\cap D=\{0\}$ and $\dim_{{\rm I\!F}_{q}}(C)+\dim_{{\rm I\!F}_{q}}(D)=nm$. Then we have following results.
@@
Let $C$ and $D$ be two ${\rm I\!F}_{q}$-linear subspaces of ${\rm I\!F}_{q^{m}}^{n}$. With respect to the inner product $B$ (see Equation (2.1)), we have the following:
@@
    $(C+D)^{\perp_{L}}=C^{\perp_{L}}\cap D^{\perp_{L}}$;
@@
    $(C+D)^{\perp_{R}}=C^{\perp_{R}}\cap D^{\perp_{R}}$;
@@
    $C^{\perp_{L}}+D^{\perp_{L}}=(C\cap D)^{\perp_{L}}$;
@@
    $C^{\perp_{R}}+D^{\perp_{R}}=(C\cap D)^{\perp_{R}}$.
@@
    Let $x\in(C+D)^{\perp_{L}}$. Then
    $\mathcal{B}(x,a)=0\text{ for all }a\in C+D\implies\mathcal{B}(x,c+d)=0\text{ for all }c\in C,d\in D.$
    If $d=0$ then $\mathcal{B}(x,c)=0~{}\text{for all}~{}c\in C\implies x\in C^{\perp_{L}}$.
    Similarly If $c=0$ then $\mathcal{B}(x,d)=0~{}\text{for all}~{}d\in D,\implies x\in D^{\perp_{L}}$.
    Hence, $x\in C^{\perp_{L}}\cap D^{\perp_{L}}.$
    For the other direction, let $y\in C^{\perp_{L}}\cap D^{\perp_{L}}$.
    Then $\mathcal{B}(y,c)=0~{}\text{for~{}all}~{}c\in C~{}\text{and}~{}\mathcal{B}(y,d)=0~{}\text{for~{}all}~{}d\in D$. That implies, $\mathcal{B}(x,c+d)=\mathcal{B}(x,c)+\mathcal{B}(x,d)=0~{}\text{for~{}all}~{}c\in C,~{}d\in D$. Hence, $x\in(C+D)^{\perp_{L}}$.
@@
    $C^{\perp_{L}}+D^{\perp_{L}}=\left((C^{\perp_{L}}+D^{\perp_{L}})^{\perp_{R}}\right)^{\perp_{L}}$ (see Equation (4)).
    Hence, $C^{\perp_{L}}+D^{\perp_{L}}=(C\cap D)^{\perp_{L}}$ (follows from 2).
@@
Let $C$ and $D$ be two ${\rm I\!F}_{q}$-linear subspaces of ${\rm I\!F}_{q^{m}}^{n}$. With respect to inner product $\mathcal{B}$ (see Equation (2.1)), the followings are equivalent.
@@
    the pair $(C,D)$ is an ACP of codes;
@@
    the pair $(C^{\perp_{L}},D^{\perp_{L}})$ is an ACP of codes;
@@
    the pair $(C^{\perp_{R}},D^{\perp_{R}})$ is an ACP of codes.
@@
In order to prove that 1)1)1 ), 2)2)2 ) and 3)3)3 ) are equivalent, it is enough to show that 1)1)1 ) and 2)2)2 ) are equivalent. Let us assume that the pair $(C,D)$ is an ACP of codes. Then $C\cap D=\{0\}$ and $\dim_{{\rm I\!F}_{q}}(C)+\dim_{{\rm I\!F}_{q}}(D)=nm$. That is, $C+_{{\rm I\!F}_{q}}D={\rm I\!F}_{q^{m}}^{n}$. Then using result in Lemma 3.1[Item 1], we have $C^{\perp_{L}}\cap D^{\perp_{L}}=\{0\}$ and from result in Equation (2), we have $\dim_{{\rm I\!F}_{q}}(C^{\perp_{L}})+\dim_{{\rm I\!F}_{q}}(D^{\perp_{L}})=nm$. Hence, $(C^{\perp_{L}},D^{\perp_{L}})$ is an ACP of codes.
Conversely, suppose that the pair $(C^{\perp_{L}},D^{\perp_{L}})$ is an ACP of codes, which gives that $C^{\perp_{L}}\cap D^{\perp_{L}}=\{0\}$ and $C^{\perp_{L}}+_{{\rm I\!F}_{q}}D^{\perp_{L}}={\rm I\!F}_{q^{m}}^{n}$. From Lemma 3.1, it follows that $C+_{{\rm I\!F}_{q}}D={\rm I\!F}_{q^{m}}^{n}$ and $C\cap D=\{0\}$. Hence, $(C,D)$ is an ACP of codes. ∎
For a matrix $G=(g_{ij})$ with entries $g_{ij}\in{\rm I\!F}_{q^{m}}$, we denote the matrix $Tr(G)=(Tr(g_{ij}))$ over ${\rm I\!F}_{q}$. The following theorem presents an characterizion of an ACP of codes.
@@
If $C$ and $D$ are two ${\rm I\!F}_{q}$-linear additive codes over ${\rm I\!F}_{q^{m}}$ of length $n$ with generator matrices $G_{1}$ and $G_{2}$, respectively. If the pair $(C,D)$ is an ACP of codes, then $rank\left(Tr\left(M\left(\begin{array}[]{cc}\pi(G_{1})\\ \pi(G_{2})\end{array}\right)P\right)\right)=n$, where $\pi$ is a field automorphism on ${\rm I\!F}_{q^{m}}$ and $\sigma~{}:~{}\{1,2,\ldots,n\}\rightarrow\{1,2,\ldots,n\}$ is a permutation with corresponding matrix $P$ such that
@@
Let assume that $rank\left(Tr\left(M\left(\begin{array}[]{cc}\pi(G_{1})\\ \pi(G_{2})\end{array}\right)P\right)\right)<n$. Then there exists $x(\neq 0)\in{\rm I\!F}_{q}^{n}$ such that $\left(Tr\left(M\left(\begin{array}[]{cc}\pi(G_{1})\\ \pi(G_{2})\end{array}\right)P\right)\right)x^{\top}=0$. This gives that $\left(Tr\left(M\left(\begin{array}[]{cc}\pi(G_{1})\\ 0\end{array}\right)P\right)\right)x^{\top}=0$ and $\left(Tr\left(M\left(\begin{array}[]{cc}0\\ \pi(G_{2})\end{array}\right)P\right)\right)x^{\top}=0$. That implies $\mathcal{B}(x,c)=0$ for all $c\in C$ and $\mathcal{B}(x,d)=0$ for all $d\in D$, that is $x\in C^{\perp_{L}}$ and $x\in D^{\perp_{L}}$. Hence, $x(\neq 0)\in C^{\perp_{L}}\cap D^{\perp_{L}}$. Then $(C^{\perp_{L}},D^{\perp_{L}})$ is not an ACP of codes and that implies $(C,D)$ is not an ACP of codes (see Theorem 3.2). This contradicts the hypothesis that $(C,D)$ is an ACP of codes. Hence $rank\left(Tr\left(M\left(\begin{array}[]{cc}\pi(G_{1})\\ \pi(G_{2})\end{array}\right)P\right)\right)=n$. ∎
We can have similar result in terms of parity check matrices as in the following Corollary.
@@
If $C$ and $D$ are two ${\rm I\!F}_{q}$-linear additive codes over ${\rm I\!F}_{q^{m}}$of length $n$ with parity check matrices $H_{1}$ and $H_{2}$, respectively. If the pair $(C,D)$ is an ACP of codes then $rank\left(Tr\left(M\left(\begin{array}[]{cc}\pi(H_{1})\\ \pi(H_{2})\end{array}\right)P\right)\right)=n$, where $\pi$ is a field automorphism on ${\rm I\!F}_{q^{m}}$ and $\sigma~{}:~{}\{1,2,\ldots,n\}\rightarrow\{1,2,\ldots,n\}$ is a permutation with corresponding matrix $P$ such that
@@
Take ${\rm I\!F}_{q}={\rm I\!F}_{2}$ and ${\rm I\!F}_{q^{m}}={\rm I\!F}_{4}$ with $\pi$ is identity automorphism on ${\rm I\!F}_{q^{m}}$. Let $C$ and $D$ be two ${\rm I\!F}_{2}$-linear additive codes over ${\rm I\!F}_{4}$ of length $3$ with generator matrices $G_{1}=\left(\begin{array}[]{ccc}1&1&0\\ \omega&\omega&0\\ \omega^{2}&0&\omega^{2}\end{array}\right)$ and $G_{1}=\left(\begin{array}[]{ccc}1&1&1\\ \omega&\omega&\omega\\ \omega&\omega&0\end{array}\right),$ respectively, where $\omega$ is a primitive root of ${\rm I\!F}_{4}$. It is easy to see that $C\cap D\neq\{0\}$, as $(\omega,~{}\omega,~{}0)\in C\cap D$. Hence $(C,D)$ is not an ACP of codes. However, $rank\left(Tr\left(\begin{array}[]{cc}G_{1}\\ G_{2}\end{array}\right)\right)=3$.
@@
Let $C$ and $D$ be two ${\rm I\!F}_{q}$-linear additive codes over ${\rm I\!F}_{q^{m}}$ of length $n$ with generator matrices $G_{1}$ and $G_{2}$ and parity check matrices $H_{1}$ and $H_{2}$, respectively. Assume that $\dim_{{\rm I\!F}_{q}}(C)+\dim_{{\rm I\!F}_{q}}(D)=mn$. Then the pair $(C,D)$ is an ACP of codes if and only if $rank\left(Tr\left(H_{2}M\left(\pi\left(G_{1}\right)P\right)^{\top}\right)\right)=rank(G_{1})$ and $rank\left(Tr\left(H_{1}M\left(\pi\left(G_{2}\right)P\right)^{\top}\right)\right)=rank(G_{2})$ where $M$ and $P$ are defined as in Theorem 3.3.
@@
Suppose that $rank\left(Tr\left(H_{2}M\left(\pi\left(G_{1}\right)P\right)^{\top}\right)\right)=rank(G_{1})$ and $rank\left(Tr\left(H_{1}M\left(\pi\left(G_{2}\right)P\right)^{\top}\right)\right)=rank(G_{2})$. To prove $(C,D)$ is an ACP of codes, we need to show that $C\cap D=\{0\}$. Let $G_{1}$ be an $k\times n$ matrix and $G_{2}$ be an $nm-k\times n$ matrix. If $x\in C\cap D$ then
$x=\alpha G_{1}~{}\text{and}~{}x=\beta G_{2}~{}\text{where}~{}\alpha\in{\rm I\!F}_{q}^{n},~{}\beta\in{\rm I\!F}_{q}^{nm-k},$
$\implies$ $\pi(\alpha)\pi(G_{1})=\pi(\beta)\pi(G_{2}),~{}\text{where}~{}\pi~{}\text{is a field automorphism on}~{}{\rm I\!F}_{q^{m}},$
$\implies$ $\pi(\alpha)\pi(G_{1})P=\pi(\beta)\pi(G_{2})P,~{}\text{where}~{}\sigma~{}:~{}\{1,2,\ldots,n\}\rightarrow\{1,2,\ldots,n\}~{}$
@@
$\text{where}~{}M=diag(\mu_{1},\mu_{2},\ldots,\mu_{n})~{}\text{is an}~{}n\times n~{}\text{matrix over}~{}{\rm I\!F}_{q^{m}}~{}\text{with}~{}\mu_{i}\in{\rm I\!F}_{q^{m}}^{*},$
@@
$\implies$ $\pi(\alpha)=0~{}\text{as}~{}rank\left(Tr\left(H_{2}M\left(\pi\left(G_{1}\right)P\right)^{\top}\right)\right)=rank(G_{1}),$
$\pi(\beta)=0~{}\text{as}~{}rank\left(Tr\left(H_{1}M\left(\pi\left(G_{2}\right)P\right)^{\top}\right)\right)=rank(G_{2}).$
Hence, $x=0$ i.e., $C\cap D=\{0\}$. As by hypothesis $\dim_{{\rm I\!F}_{q}}(C)+\dim_{{\rm I\!F}_{q}}(D)=mn$, the pair $(C,D)$ is an ACP of codes.
Conversely, let us suppose that the pair $(C,D)$ is an ACP of codes. Let $rank\left(Tr\left(H_{2}M\left(\pi\left(G_{1}\right)P\right)^{\top}\right)\right)<rank(G_{1})$. Then there exists $x(\neq 0)\in{\rm I\!F}_{q}^{n}$ such that $\left(Tr\left(H_{2}M\left(\pi\left(G_{1}\right)P\right)^{\top}\right)\right)x^{\top}=0$. That implies, $\left(Tr\left(H_{2}M\left(\pi\left(xG_{1}\right)P\right)^{\top}\right)\right)=0$. That is, $xG_{1}\in D$. As already, $xG_{1}\in C$, $xG_{1}(\neq 0)\in C\cap D$, which contradicts the fact that the $(C,D)$ is an ACP of codes. Therefore, $rank\left(Tr\left(H_{2}D\pi\left(G_{1}\right)P\right)^{\top}\right)=rank(G_{1})$. Similarly, we can prove that $rank\left(Tr\left(H_{1}M\left(\pi\left(G_{2}\right)P\right)^{\top}\right)\right)=rank(G_{2})$. ∎

### v3

@@ 3 Characterization of ACP of codes
For two ${\rm I\!F}_{q}$-linear additive codes over ${\rm I\!F}_{q^{m}}$, $\mathcal{C}$ and $\mathcal{D}$, the pair $(\mathcal{C},\mathcal{D})$ is called an additive complementary pair (ACP) of codes if $\mathcal{C}\oplus_{{\rm I\!F}_{q}}\mathcal{D}={\rm I\!F}_{q^{m}}^{n}$, i.e., $\mathcal{C}\cap\mathcal{D}=\{0\}$ and $\mathcal{C}+_{{\rm I\!F}_{q}}\mathcal{D}=\{\alpha\mathbf{c}+\beta\mathbf{d}|\alpha,\beta\in{\rm I\!F}_{q},\mathbf{c}\in\mathcal{C},\mathbf{d}\in\mathcal{D}\}={\rm I\!F}_{q^{m}}^{n}$. Equivalently, a pair $(\mathcal{C},\mathcal{D})$ is an ACP of codes if and only if $\mathcal{C}\cap\mathcal{D}=\{0\}$ and $\dim_{{\rm I\!F}_{q}}(\mathcal{C})+\dim_{{\rm I\!F}_{q}}(\mathcal{D})=nm$. Then, we have the following results.
@@
Let $\mathcal{C}$ and $\mathcal{D}$ be two ${\rm I\!F}_{q}$-linear subspaces of ${\rm I\!F}_{q^{m}}^{n}$. With respect to the inner product $B$ (see Equation (2.1)), we have
@@
    $(\mathcal{C}+\mathcal{D})^{\perp_{L}}=\mathcal{C}^{\perp_{L}}\cap\mathcal{D}^{\perp_{L}}$;
@@
    $(\mathcal{C}+\mathcal{D})^{\perp_{R}}=\mathcal{C}^{\perp_{R}}\cap\mathcal{D}^{\perp_{R}}$;
@@
    $\mathcal{C}^{\perp_{L}}+\mathcal{D}^{\perp_{L}}=(\mathcal{C}\cap\mathcal{D})^{\perp_{L}}$;
@@
    $\mathcal{C}^{\perp_{R}}+\mathcal{D}^{\perp_{R}}=(\mathcal{C}\cap\mathcal{D})^{\perp_{R}}$.
@@
[…]
@@
    As $\mathcal{C}^{\perp_{L}}+\mathcal{D}^{\perp_{L}}=\left((\mathcal{C}^{\perp_{L}}+\mathcal{D}^{\perp_{L}})^{\perp_{R}}\right)^{\perp_{L}}$ (see Equation (4)), $\mathcal{C}^{\perp_{L}}+\mathcal{D}^{\perp_{L}}=(\mathcal{C}\cap\mathcal{D})^{\perp_{L}}$ (follows from 2).
@@
Let $\mathcal{C}$ and $\mathcal{D}$ be two ${\rm I\!F}_{q}$-linear subspaces of ${\rm I\!F}_{q^{m}}^{n}$. With respect to the inner product $\mathcal{B}$ (see Equation (2.1)), the following are equivalent.
@@
    the pair $(\mathcal{C},\mathcal{D})$ is an ACP of codes;
@@
    the pair $(\mathcal{C}^{\perp_{L}},\mathcal{D}^{\perp_{L}})$ is an ACP of codes;
@@
    the pair $(\mathcal{C}^{\perp_{R}},\mathcal{D}^{\perp_{R}})$ is an ACP of codes.
@@
In order to prove that 1)1)1 ), 2)2)2 ) and 3)3)3 ) are equivalent, it is enough to show that 1)1)1 ) and 2)2)2 ) are equivalent. Let us assume that the pair $(\mathcal{C},\mathcal{D})$ is an ACP of codes. Then $\mathcal{C}\cap\mathcal{D}=\{0\}$ and $\dim_{{\rm I\!F}_{q}}(\mathcal{C})+\dim_{{\rm I\!F}_{q}}(\mathcal{D})=nm$. That is, $\mathcal{C}+_{{\rm I\!F}_{q}}\mathcal{D}={\rm I\!F}_{q^{m}}^{n}$. Then using result in 1)1)1 ) of Lemma 3.1, we have $\mathcal{C}^{\perp_{L}}\cap\mathcal{D}^{\perp_{L}}=\{0\}$ and since $\dim_{{\rm I\!F}_{q}}(\mathcal{C})+\dim_{{\rm I\!F}_{q}}(\mathcal{D})=nm$ from result in Equation (2), we have $\dim_{{\rm I\!F}_{q}}(\mathcal{C}^{\perp_{L}})+\dim_{{\rm I\!F}_{q}}(\mathcal{D}^{\perp_{L}})=nm$. Hence, $(\mathcal{C}^{\perp_{L}},\mathcal{D}^{\perp_{L}})$ is an ACP of codes.
Conversely, suppose that the pair $(\mathcal{C}^{\perp_{L}},\mathcal{D}^{\perp_{L}})$ is an ACP of codes, which gives that $\mathcal{C}^{\perp_{L}}\cap\mathcal{D}^{\perp_{L}}=\{0\}$ and $\mathcal{C}^{\perp_{L}}+_{{\rm I\!F}_{q}}\mathcal{D}^{\perp_{L}}={\rm I\!F}_{q^{m}}^{n}$. From Lemma 3.1, it follows that $\mathcal{C}+_{{\rm I\!F}_{q}}\mathcal{D}={\rm I\!F}_{q^{m}}^{n}$ and $\mathcal{C}\cap\mathcal{D}=\{0\}$. Hence, $(\mathcal{C},\mathcal{D})$ is an ACP of codes. ∎
For a matrix $G=(g_{ij})$ with entries $g_{ij}\in{\rm I\!F}_{q^{m}}$, we denote the matrix $Tr(G)=(Tr(g_{ij}))$ over ${\rm I\!F}_{q}$. The following theorem presents a characterization of an ACP of codes.
@@
If $\mathcal{C}$ and $\mathcal{D}$ are two ${\rm I\!F}_{q}$-linear additive codes over ${\rm I\!F}_{q^{m}}$ of length $n$ with generator matrices $G_{1}$ and $G_{2}$, respectively. If the pair $(\mathcal{C},\mathcal{D})$ is an ACP of codes, then $rank\left(Tr\left(\begin{array}[]{cc}M\left(\pi(G_{1})P\right)^{\top}\\ M\left(\pi(G_{2})P\right)^{\top}\end{array}\right)\right)=n$, where $\pi$ is a field automorphism on ${\rm I\!F}_{q^{m}}$ and $\sigma~{}:~{}\{1,2,\ldots,n\}\rightarrow\{1,2,\ldots,n\}$ is a permutation with corresponding matrix $P$ such that
@@
Let assume that $rank\left(Tr\left(\begin{array}[]{cc}M\left(\pi(G_{1})P\right)^{\top}\\ M\left(\pi(G_{2})P\right)^{\top}\end{array}\right)\right)<n$. Then there exists nonzero $\mathbf{x}\in{\rm I\!F}_{q}^{n}$ such that $\left(Tr\left(\begin{array}[]{cc}M\left(\pi(G_{1})P\right)^{\top}\\ M\left(\pi(G_{2})P\right)^{\top}\end{array}\right)\right)\mathbf{x}^{\top}=0$. Then $\left(Tr\left(\begin{array}[]{cc}M\left(\pi(G_{1})P\right)^{\top}\\ \end{array}\right)\right)\mathbf{x}^{\top}=\left(Tr\left(\begin{array}[]{cc}M\left(\pi(G_{2})P\right)^{\top}\end{array}\right)\right)\mathbf{x}^{\top}=0$. That implies $\mathcal{B}(\mathbf{x},\mathbf{c})=0$ for all $\mathbf{c}\in\mathcal{C}$ and $\mathcal{B}(\mathbf{x},\mathbf{d})=0$ for all $\mathbf{d}\in\mathcal{D}$, that is $\mathbf{x}\in\mathcal{C}^{\perp_{L}}$ and $\mathbf{x}\in\mathcal{D}^{\perp_{L}}$. Hence, $\mathbf{x}(\neq 0)\in\mathcal{C}^{\perp_{L}}\cap\mathcal{D}^{\perp_{L}}$. Then $(\mathcal{C}^{\perp_{L}},\mathcal{D}^{\perp_{L}})$ is not an ACP of codes and that implies $(\mathcal{C},\mathcal{D})$ is not an ACP of codes (see Theorem 3.2). This contradicts the hypothesis that $(\mathcal{C},\mathcal{D})$ is an ACP of codes. Hence $rank\left(Tr\left(\begin{array}[]{cc}M\left(\pi(G_{1})P\right)^{\top}\\ M\left(\pi(G_{2})P\right)^{\top}\\ \end{array}\right)\right)=n$. ∎
Example 3.4.
Take ${\rm I\!F}_{q}={\rm I\!F}_{2}$ and ${\rm I\!F}_{q^{m}}={\rm I\!F}_{4}$ with $\pi$ is the automorphism on ${\rm I\!F}_{q^{m}}$, define as $\pi(x)=x^{2}$. Let $\mathcal{C}$ and $\mathcal{D}$ be two ${\rm I\!F}_{2}$-linear additive codes over ${\rm I\!F}_{4}$ of length $6$ with generator matrices
$G_{1}=\left(\begin{array}[]{cccccc}1&0&0&0&0&1\\ \omega&0&0&0&0&\omega\\ 0&1&0&0&0&1\\ 0&\omega&0&0&0&\omega\\ 0&0&1&0&0&1\\ 0&0&\omega&0&0&\omega\\ 0&0&0&1&0&1\\ 0&0&0&\omega&0&\omega\\ \end{array}\right)\text{ and }G_{2}=\left(\begin{array}[]{cccccc}0&0&0&0&1&0\\ 0&0&0&0&\omega&0\\ 1&1&1&1&0&1\\ \omega&\omega&\omega&\omega&0&\omega\\ \end{array}\right),$
respectively, where $\omega$ is a primitive root of ${\rm I\!F}_{4}$. It is easy to see that $(\mathcal{C},\mathcal{D})$ is an ACP of codes with respect to $P$ such that
$P_{ij}=\left\{\begin{array}[]{ll}1&\textrm{if~{} off-diagonal element };\\ 0&\textrm{otherwise,}\end{array}\right.$
$M=diag(\mu_{1},\mu_{2},\ldots,\mu_{n})$ is an $n\times n$ matrix over ${\rm I\!F}_{4}$ with $\mu_{i}\in{\rm I\!F}_{4}^{*}$. However, $rank\left(Tr\left(\begin{array}[]{cc}M\left(\pi(G_{1})P\right)^{\top}\\ M\left(\pi(G_{2})P\right)^{\top}\end{array}\right)\right)=6$, which is coincide with Theorem 3.3.
We can have similar results in terms of parity check matrices as in the following Corollary.
@@
If $\mathcal{C}$ and $\mathcal{D}$ are two ${\rm I\!F}_{q}$-linear additive codes over ${\rm I\!F}_{q^{m}}$of length $n$ with parity check matrices $H_{1}$ and $H_{2}$, respectively. If the pair $(\mathcal{C},\mathcal{D})$ is an ACP of codes then $rank\left(Tr\left(\begin{array}[]{cc}M\left(\pi(H_{1})P\right)^{\top}\\ M\left(\pi(H_{2})P\right)^{\top}\end{array}\right)\right)=n$, where $\pi$ is a field automorphism on ${\rm I\!F}_{q^{m}}$ and $\sigma~{}:~{}\{1,2,\ldots,n\}\rightarrow\{1,2,\ldots,n\}$ is a permutation with corresponding matrix $P$ such that
@@
Take ${\rm I\!F}_{q}={\rm I\!F}_{2}$ and ${\rm I\!F}_{q^{m}}={\rm I\!F}_{4}$ with $\pi$ is identity automorphism on ${\rm I\!F}_{q^{m}}$. Let $\mathcal{C}$ and $\mathcal{D}$ be two ${\rm I\!F}_{2}$-linear additive codes over ${\rm I\!F}_{4}$ of length $3$ with generator matrices
$G_{1}=\left(\begin{array}[]{ccc}1&1&0\\ \omega&\omega&0\\ \omega^{2}&0&\omega^{2}\end{array}\right)\text{ and }G_{1}=\left(\begin{array}[]{ccc}1&1&1\\ \omega&\omega&\omega\\ \omega&\omega&0\end{array}\right),$
respectively, where $\omega$ is a primitive root of ${\rm I\!F}_{4}$. It is easy to see that $C\cap D\neq\{0\}$, as $(\omega,~{}\omega,~{}0)\in C\cap D$. Hence $(\mathcal{C},\mathcal{D})$ is not an ACP of codes. However, $rank\left(Tr\left(\begin{array}[]{cc}G_{1}\\ G_{2}\end{array}\right)\right)=3$.
@@
Let $\mathcal{C}$ and $\mathcal{D}$ be two ${\rm I\!F}_{q}$-linear additive codes over ${\rm I\!F}_{q^{m}}$ of length $n$ with generator matrices $G_{1}$ and $G_{2}$ and parity check matrices $H_{1}$ and $H_{2}$, respectively. Assume that $\dim_{{\rm I\!F}_{q}}(\mathcal{C})+\dim_{{\rm I\!F}_{q}}(\mathcal{D})=mn$. Then the pair $(\mathcal{C},\mathcal{D})$ is an ACP of codes if and only if $rank\left(Tr\left(H_{2}M\left(\pi\left(G_{1}\right)P\right)^{\top}\right)\right)=rank(G_{1})$ and $rank\left(Tr\left(H_{1}M\left(\pi\left(G_{2}\right)P\right)^{\top}\right)\right)=rank(G_{2})$ where $M$ and $P$ are defined as in Theorem 3.3.
@@
Suppose that $rank\left(Tr\left(H_{2}M\left(\pi\left(G_{1}\right)P\right)^{\top}\right)\right)=rank(G_{1})$ and $rank\left(Tr\left(H_{1}M\left(\pi\left(G_{2}\right)P\right)^{\top}\right)\right)=rank(G_{2})$. To prove $(\mathcal{C},\mathcal{D})$ is an ACP of codes, we need to show that $\mathcal{C}\cap\mathcal{D}=\{0\}$. Let $G_{1}$ be an $k\times n$ matrix and $G_{2}$ be an $nm-k\times n$ matrix. If $\mathbf{x}\in\mathcal{C}\cap\mathcal{D}$ then
$\mathbf{x}=\mathbf{\alpha}G_{1}~{}\text{and}~{}\mathbf{x}=\mathbf{\beta}G_{2}~{}\text{where}~{}\mathbf{\alpha}\in{\rm I\!F}_{q}^{n},~{}\mathbf{\beta}\in{\rm I\!F}_{q}^{nm-k},$
$\implies$ $\pi(\mathbf{\alpha})\pi(G_{1})=\pi(\mathbf{\beta})\pi(G_{2}),~{}\text{where}~{}\pi~{}\text{is a field automorphism on}~{}{\rm I\!F}_{q^{m}},$
$\implies$ $\pi(\mathbf{\alpha})\pi(G_{1})P=\pi(\mathbf{\beta})\pi(G_{2})P,~{}\text{where}~{}\sigma~{}:~{}\{1,2,\ldots,n\}\rightarrow\{1,2,\ldots,n\}~{}$
@@
where $\text{where }M=diag(\mu_{1},\mu_{2},\ldots,\mu_{n})\text{ is an }n\times n\text{ matrix over }{\rm I\!F}_{q^{m}}\text{ with }\mu_{i}\in{\rm I\!F}_{q^{m}}^{*},$
@@
$\implies$ $\pi(\alpha)=0\text{ as }rank\left(Tr\left(H_{2}M\left(\pi\left(G_{1}\right)P\right)^{\top}\right)\right)=rank(G_{1}),$
$\pi(\beta)=0\text{ as }rank\left(Tr\left(H_{1}M\left(\pi\left(G_{2}\right)P\right)^{\top}\right)\right)=rank(G_{2})$
Hence, $\mathbf{x}=0$ i.e., $\mathcal{C}\cap\mathcal{D}=\{0\}$. As by hypothesis $\dim_{{\rm I\!F}_{q}}(\mathcal{C})+\dim_{{\rm I\!F}_{q}}(\mathcal{D})=mn$, the pair $(\mathcal{C},\mathcal{D})$ is an ACP of codes.
Conversely, let us suppose that the pair $(\mathcal{C},\mathcal{D})$ is an ACP of codes. Let $rank\left(Tr\left(H_{2}M\left(\pi\left(G_{1}\right)P\right)^{\top}\right)\right)<rank(G_{1})$. Then there exists $\mathbf{x}(\neq 0)\in{\rm I\!F}_{q}^{n}$ such that $\left(Tr\left(H_{2}M\left(\pi\left(G_{1}\right)P\right)^{\top}\right)\right)\mathbf{x}^{\top}=0$. That implies, $\left(Tr\left(H_{2}M\left(\pi\left(\mathbf{x}G_{1}\right)P\right)^{\top}\right)\right)=0$. That is, $\mathbf{x}G_{1}\in\mathcal{D}$. As already, $\mathbf{x}G_{1}\in\mathcal{C}$, $\mathbf{x}G_{1}(\neq 0)\in\mathcal{C}\cap\mathcal{D}$, which contradicts the fact that the $(\mathcal{C},\mathcal{D})$ is an ACP of codes. Therefore, $rank\left(Tr\left(H_{2}D\pi\left(G_{1}\right)P\right)^{\top}\right)=rank(G_{1})$. Similarly, we can prove that $rank\left(Tr\left(H_{1}M\left(\pi\left(G_{2}\right)P\right)^{\top}\right)\right)=rank(G_{2})$. ∎